# Data Taggants: Dataset Ownership Verification via Harmless Targeted Data Poisoning

### v1

@@ 3.2 Signing the data
To prevent from negatively impacting model performances, we choose key images $x^{(key)}_{i}$ to be out-of-distribution data points. More precisely, we generate key images by sampling each pixel value uniformly and sample key labels $y^{(key)}_{i}$ uniformly at random in $\mathcal{Y}$.
@@
We evenly split the signing set into K parts $(\mathcal{D}_{S_{i}})_{i=1}^{K}$, where $\mathcal{D}_{S_{i}}$ is associated to key $i$. Following clean-label data poisoning (Geiping et al., 2020), each sample $(x_{j},y_{j})$ in $\mathcal{D}_{S_{i}}$ satisfies $y_{j}=y^{(key)}_{i}$. With this split, the taggant optimization problem, described later, is decomposable into one problem for each key, which we solve in parallel.
Figure 2: Illustration of our method: Alice optimizes image-wise signatures for images in the signing set by maximizing the alignment between the gradients of the signed images $\nabla_{\theta}^{(signed)}$, and the gradient of the key $\nabla_{\theta}^{(key)}$. The resulting images and their labels are the data taggant.
@@
We approximate $\mathbb{E}_{t}\left[\nabla_{\theta}\mathcal{L}^{A}_{\theta^{*}}(t(x_{j}+\delta_{j}),y_{j})\right]$ by repeating each sample $R$ times, applying $R$ randomly sampled data augmentations $t_{r},r\in\{1,\ldots,R\}$, and averaging the resulting gradients. We use differentiable versions of common data augmentations to optimize through them. The taggant objective function $\mathcal{T}(\Delta)$ is illustrated in Figure 2, computes the perturbation by aligning the gradients of the keys with the gradient of perturbed images. More precisely, $\mathcal{T}(\Delta)=\sum_{i=1}^{K}\mathcal{T}_{i}$ with:
$\mathcal{T}_{i}(\Delta)=-\cos\Big{(}\nabla_{\theta}\mathcal{L}^{A}_{\theta}(x^{(key)}_{i},y^{(key)}_{i}),\sum\limits_{j\in\mathcal{D}_{S_{i}}}\frac{1}{R}\sum\limits_{r=1}^{R}\nabla_{\theta}\mathcal{L}^{A}_{\theta}(t_{r}(x_{j}+\delta_{j}),y_{j})\Big{)}$ (2)
@@
To improve the stealthiness of our approach, we add a term of differentiable perceptual loss $\mathcal{L}_{perc}$ to the taggant function, using the LPIPS metric Zhang et al. (2018), which relies on the visual features extracted by a VGG model. Given a weight $\lambda$ for $\mathcal{L}_{perc}$, the final optimization problem for taggants is $\min\limits_{\Delta\in\mathcal{C}_{\mathcal{S}}}\frac{1}{K}\sum_{i=1}^{K}\mathcal{T}_{i}(\Delta)+\lambda\mathcal{L}_{perc}(\Delta)$.
@@
Since the problem is non-convex, the algorithm may be trapped in local minima. In practice, we observed that using multiple random restarts of initial signatures for each key improved performance. We chose the best according to the taggant objective function.
@@ 3.3 Detection
The detection phase ( 2 in Figure 1) relies on checking a suspicious model’s predictions on the secret keys. In order to control for false positives, we determine the key accuracy of an honest model.
@@
Our null hypothesis is $\mathcal{H}_{0}$: “model $h$ was not trained on the signed dataset”. Given the set of keys $\mathcal{D}_{K}$ where the key labels are chosen uniformly at random, under the null hypothesis, model $h$ cannot have better-than-chance probability of predicting $y^{(key)}_{i}$ given $x^{(key)}_{i}$.
[…]
@@ 4 Experiments
In this section, we empirically evaluate the effectiveness, stealthiness, and robustness of data taggants.
@@
We use ImageNet1k (Deng et al., 2009) with Vision Transformers (DeIT) and Residual Networks (ResNet) models with different sizes and state-of-the-art training recipes from Wightman et al. (2021) and Touvron et al. (2022) (see Appendix A.2 for details). When generating taggants, the signing sets’ sizes use a budget $B$ of the total dataset size, $S=B\times N$, with $B=10^{-3}$ unless stated otherwise. We generate 20 secret keys, using 8 random restart per key and keep the 10 best keys according to the taggant objective function. The weight of the perceptual loss is fixed to $\lambda=0.01$. These parameters were chosen in preliminary experiments and fixed afterwards. $\varepsilon$ in the constraint set $\mathcal{C}$ is fixed to $16/255$, a common value in the data poisoning literature.
In each experiment, we train one model for Alice and run the detection test on 4 repetitions of Bob’s training 4 using different random initializations to compute standard deviations. Similarly to Sablayrolles et al. (2020), we combine the p-values of the 4 tests with Fisher’s method Fisher (1970). The result is denoted $\log_{10}p$ and is commensurable to the base-10 logarithm of a p-value. Given the large number of experiments, we train all models for 100 epochs to save compute, even though the best performances for DeIT are achieved after 800 epochs. As sanity check, we confirm with a run of 800 epochs that the effectiveness results hold. The results are in Table 9 in Appendix A.1.
Table 1: Validation accuracy and detection performance when both Alice and Bob use DeIT-small models with the three-augment data augmentation and various budgets.
Method Budget $B$ Validation accuracy Top-$k$ keys accuracy PSNR
@@
Clean training $0.0\%$ $\mathbf{64.2\pm 0.4}$
$0.01\%$ $\mathbf{64.2\pm 0.3}$ $15.0\pm 10.0$ (-9.2) $\mathbf{100.0\pm 0.0}$ (-74.0) -
$0.1\%$ $63.8\pm 1.1$ $85.0\pm 19.1$ (-91.7) $\mathbf{100.0\pm 0.0}$ (-74.0) -
$0.01\%$ $63.4\pm 0.7$ $7.5\pm 5.0$ (-3.3) $10.0\pm 0.0$ ( -1.8) $20.0$
$0.1\%$ $63.6\pm 0.6$ $10.0\pm 0.0$ (-4.9) $\mathbf{55.0\pm 5.8}$ (-29.7) $20.0$
Data taggants (Our method) $0.001\%$ $63.7\pm 0.9$ $0.0\pm 0$ 0.0 (0.0) $0.0\pm 0$ 0.0 (0.0) $26.6$
$0.01\%$ $63.6\pm 0.9$ $0.0\pm 0$ 0.0 (0.0) $5.0\pm 5.8$ (-0.5) $27.3$
$0.1\%$ $\mathbf{64.2\pm 0.6}$ $10.0\pm 0.0$ (-4.9) $\mathbf{87.5\pm 5.0}$ (-59.6) $\mathbf{27.9}$
@@
Key source Budget $B$ Validation accuracy Top-$k$ keys accuracy PSNR
@@
None $0.0\%$ $64.2\pm 0.4$
Test images $0.001\%$ $63.7\pm 0.4$ $0.0\pm 0$ 0.0 (0.0) $7.5\pm 5.0$ (-1.1) $27.5$
$0.01\%$ $63.4\pm 1.0$ $0.0\pm 0$ 0.0 (0.0) $0.0\pm 0$ 0.0 ( 0.0) $27.9$
$0.1\%$ $63.9\pm 0.6$ $0.0\pm 0$ 0.0 (0.0) $27.5\pm 5.0$ (-10.4) $28.4$
Random (Our method) $0.001\%$ $63.7\pm 0.9$ $0.0\pm 0$ 0.0 (0.0) $0.0\pm 0$ 0.0 ( 0.0) $26.6$
$0.01\%$ $63.6\pm 0.9$ $0.0\pm 0$ 0.0 (0.0) $5.0\pm 5.8$ (-0.5) $27.3$
@@ 4.1 Effectiveness
We first compare the effectiveness of data taggants to two baselines. First, “naive canary”, where copies of the private keys are added into the training set. This serves as a topline in terms of detection performance, but it is not viable as tracing mechanism since it is not stealth at all. Second, “transparency”, where we linearly interpolate between the keys and images of the signing set with a weight $\gamma=0.2$, $x^{\prime}=\gamma x^{(key)}+(1-\gamma)x$. This value for $\gamma$ was chosen by visual inspection, as a small value for which the key is still visible. In this comparison, we use DeIT-small models with the three-augment (3A) data augmentation (Touvron et al., 2022) for Alice and Bob, with three different budgets $B$, $0.1\%$, $0.01\%$ and $0.001\%$.
The results are given in Table 1 (more results can be found in Table 7 in Appendix). First, we observe that all methods have roughly the same validation acuracy as a model trained on clean data (differences within noise). In particular, this suggests that our taggants do not negatively impact model performances at these budget levels. In terms of detection, Naive canary works best as expected, with perfect key top-10 accuracy for $B\geq 0.01\%$. While Transparency works better than data taggants for smaller budgets, data taggants has much higher top-10 accuracy for $B=0.1\%$ ($87.5\%$ vs $55.0\%$). In addition, the PSNR (signal to noise ratio) is lower for transparency, which suggests that with a budget as small as $0.1\%$, data taggants already dominate the transparency baseline in this experimental setting, and are able to detect dishonest models with very high confidence ($\log_{10}p=-59.6$).
We perform an additional comparison with another baseline where test data points are used as keys rather than random patterns, which is similar to repurposing a data poisoning attack to dataset ownership verification. Table 2 shows the results. With comparable PSNR, using random keys leads to much better detection accuracy and confidence for $B=0.1\%$ than using test data, justifying our design choice.
@@ 4.2 Stealthiness
The stealthiness of data taggants is essential but difficult to measure beyond PSNR. As a best effort, we address the scenario where Bob tries to locate the signed data using either visual inspection, or, following (Tang et al., 2023), using anomaly detection algorithms. Similarly to the previous section, we use taggants crafted by DeIT-small using 3A augmentation, with a budget $B=0.1\%$.
@@
We provide examples of taggants crafted with and without perceptual loss in Figure 3 (more examples are given in Figure 5, 6 and 7 in Appendix A.1). We observe that while most of them appear unaltered, some of them display visible patterns of weak intensity that could as well come from natural film grain. Although with PSNR ¡ 30 dB (see Table 1) which is considered low according to image processing standards, we believe that the data taggants are reasonably stealth against visual inspection, particularly when Bob has to find them among a whole dataset.
@@
Most methods for out-of-distribution (OOD) detection (Chen et al., 2020) rely on training a model on a set of in-lier data to then be able to detect outliers. Here, even though in reality Bob does not have access to the original clean data, we consider a topline detection scenario where Bob has access to a “clean” model, i.e., a model trained on Alice’s original dataset and knows which classes contain data taggants when using OOD detection algorithms.
We consider two off-the-shelf OOD detection algorithms, the first one using k-NN (Kuan & Mueller, 2022) from the cleanlab project [^2] https://github.com/cleanlab/cleanlab, and the clustering method DBSCAN (Schubert et al., 2017) from the scikit-learn project [^3] https://scikit-learn.org/stable/index.html. Both methods use the features computed from the clean model.
The results are given in Figure 4. We observe that neither method detect data taggants – less than $1.1\%$ taggant detection rate for the best setting of $k$ in k-NN, when considering the top-1000 samples with highest outlier scores. Interestingly, when computing DBSCAN for different detection thresholds (Figure 4 (b)), we observe that DBSCAN exhibit performances that are significantly worse than random. Our explanation is that DBSCAN computes clusters and select outliers as isolated points or clusters. By manually analyzing the clusters, we observe that some of them contain a lot of data taggants, most likely because their embeddings are lumped together and do not appear as anomalies.
@@
Data poisoning detection.
Since data taggants rely on a data poisoning mechanism, we suggest to use data poisoning detection methods to detect data taggants. Table 3 shows the percentage of data taggants detected as outliers by three data poisoning detection methods: Deep k-NN, Activation Clustering and Spectral Clustering. The results show that data taggants cannot be detected by these methods, which suggests for strong stealthiness of data taggants.
Table 3: Percentage of total data and data taggants detected as outliers by data poisoning detection methods.
Method % of total data detected % of data taggants detected
Deep k-NN $7.7$ $6.9$
Activation Clustering $51.2$ $48.7$
Spectral Clustering $10.8$ $12.7$
(a) k-NN
(b) DBSCAN
Figure 4: Outlier detection based on the features of the signed dataset $\tilde{\mathcal{D}}_{A}$

### v2

@@ 3.2 Signing the data in practice
To prevent from negatively impacting model performance, we choose key images $x^{(key)}_{i}$ to be out-of-distribution data points. More precisely, we generate key images by sampling each pixel value uniformly and sample key labels $y^{(key)}_{i}$ uniformly at random in $\mathcal{Y}$. Since no natural behavior is expected from the model on these keys, enforcing a specific behavior on them should not induce particular errors, as opposed to backdoor watermarking approaches.
@@
We evenly split the signing set into K parts $(\mathcal{D}_{S_{i}})_{i=1}^{K}$, where $\mathcal{D}_{S_{i}}$ is associated to key $i$. Following clean-label data poisoning (Geiping et al., 2020), each sample $(x_{j},y_{j})$ in $\mathcal{D}_{S_{i}}$ satisfies $y_{j}=y^{(key)}_{i}$. While Witches’ Brew in the multi-targets case performs gradient matching to align the poisonous gradients with the averaged targeted gradients, we instead suggest to solve $K$ gradient matching problems between each part $\mathcal{D}_{S_{i}}$ and the associated key $(x^{(key)}_{i},y^{(key)}_{i})$. The taggants optimization problem, described below, is thus decomposable into one problem for each key, which blue can be solved independently in parallel.
Figure 2: Illustration of our method: Alice optimizes image-wise signatures for images in the signing set by maximizing the alignment between the gradients of the signed images $\nabla_{\theta}^{(signed)}$, and the gradient of the key $\nabla_{\theta}^{(key)}$. The resulting images and their labels are the data taggants.
@@
We use differentiable versions of data augmentations to optimize through them, similarly to Witches’ Brew but increasing the set of considered data augmentations (see Table 13 in Appendix). At each gradient matching optimization step, we approximate $\mathbb{E}_{t}\left[\nabla_{\theta}\mathcal{L}^{A}_{\theta^{*}}(t(x^{(sign)}_{j}+\delta_{j}),y^{(sign)}_{j})\right$] by resampling and applying $R$ randomly sampled data augmentations $t_{r},r\in\{1,\ldots,R\}$ (while $R=1$ in Witches’ Brew), and averaging the resulting gradients. The taggant objective function $\mathcal{T}(\Delta)$, illustrated in Figure 2, computes the alignment of perturbed images’ gradients with the keys’ gradients. More precisely, $\mathcal{T}(\Delta)=\sum_{i=1}^{K}\mathcal{T}_{i}/K$ with:
[…]
@@
To improve the stealthiness of our approach, we introduce the use of a differentiable perceptual loss term $\mathcal{L}_{perc}$ to the taggant function, using the LPIPS metric (Zhang et al., 2018), which relies on the visual features extracted by a VGG model. Given a weight $\lambda$ for $\mathcal{L}_{perc}$, the final optimization problem for taggants is $\min\limits_{\Delta\in\mathcal{C}_{\mathcal{S}}}\frac{1}{K}\sum_{i=1}^{K}\mathcal{T}_{i}(\Delta)+\lambda\mathcal{L}_{perc}(\Delta)$.
@@
Since the problem to solve is non-convex, the algorithm may be trapped in local minima. In practice, and similarly to Witches’ Brew, we observed that using multiple random restarts of initial signatures for each key improved performance. We chose, for each key, the best crafted taggants among all restarts according to the taggant objective function.
@@ 3.3 Detection
The detection phase ( 2 in Figure 1) relies on checking a suspicious model’s predictions on the secret keys. We consider the detection to be successful when a model displays a top-$k$ accuracy on the set of keys to be at least a threshold $0\leq\tau\leq 1$. In order to control for false positives, we determine the top-$k$ accuracy of an honest model on the set of keys and set $\tau$ to be higher. The choice of $k$ and $\tau$ balances between the false positive rate (FPR, wrongfully detecting a benign model) and the true positive rate (TPR, correctly detecting a model trained on $\tilde{\mathcal{D}}_{A}$, i.e. detection rate). Given Bob’s model’s top-$k$ accuracy on the set of keys, Alice can derive a corresponding $p$-value for observing similar accuracy on a benign model. If that $p$-value is deemed low enough, it can be concluded that Bob’s model was trained on Alice’s dataset.
@@
Our null hypothesis is $\mathcal{H}_{0}$: “model $h$ was not trained on the signed dataset”.
Proposition 1.
Under $\mathcal{H}_{0}$, model $h$ has, in expectation, a top-$k$ accuracy of $\frac{k}{|\mathcal{Y}|}$ on the set of keys $\mathcal{D}_{K}$, where $|\mathcal{Y}|$ is the number of possible labels.
Proof.
Under $\mathcal{H}_{0}$, for a given key, correctly predicting the random label $y^{(key)}$ based on a top-$k$ prediction given $x^{(key)}$ can be modeled by a random variable following a Bernoulli distribution of parameter ${k}/{|\mathcal{Y}|}$. Since all the labels are independently drawn, the number of correct predictions on the set of keys follows a binomial distribution with parameters $(K,{k}/{|\mathcal{Y}|})$ and an expectation of ${K\times k}/{|\mathcal{Y}|}$. Hence the expected accuracy over $K$ keys is ${k}/{|\mathcal{Y}|}$. ∎
Statistical test.
Let us denote by $\text{top-}k(h,\mathcal{D}_{K})$ the number of keys for which the label $y^{(key)}_{i}$ is in the top-$k$ predictions $h(x^{(key)}_{i})$ for $x^{(key)}_{i}\in\mathcal{D}_{K}$. Under the null hypothesis, Proposition 1 gives that $\text{top-}k(h,\mathcal{D}_{K})$ follows a binomial distribution with parameters $(K,k/|\mathcal{Y}|$) and can be subject to a binomial test. The p-value of the binomial test is then given by the tail of a random variable $Z_{k}$ following a binomial distribution $Bin(K,\frac{k}{|\mathcal{Y}|})$:
$p=\mathbb{P}(Z_{k}\geq\text{top-}k(h,\mathcal{D}_{K}))=\sum_{z=\text{top-}k(h,\mathcal{D}_{K})}^{K}\binom{K}{z}\Big(\frac{k}{|\mathcal{Y}|}\Big)^{z}\Big(\frac{|\mathcal{Y}|-k}{|\mathcal{Y}|}\Big)^{K-z}$
@@
Previous works on DOV have proposed to use a pair-wise t-test (Maini, 2021; Li et al., 2022; Wenger et al., 2022; Li et al., 2023; Tang et al., 2023; Guo et al., 2023) or a Wilcoxon signed-rank test (Li et al., 2023; Tang et al., 2023) on the model’s predictions. However, the hypotheses these works test for rely (without any theoretical grounding) on the assumption that a benign model must display, on average, similar predictions on clean images and verification images, may they be watermarked images (Li et al., 2022; Wenger et al., 2022; Li et al., 2023; Tang et al., 2023) or particular private images (Maini, 2021; Guo et al., 2023). Our detection scheme, on the other hand, do not need any assumption as it allows us to exactly characterize the behavior of a benign model on the keys.
@@ 4 Experiments
In this section, we empirically evaluate the effectiveness, stealthiness, and robustness of data taggants. Experiments are run on a realistic setting to assess the practicality of our approach.
@@
We use ImageNet1k (Deng et al., 2009) with Vision Transformers (DeIT) and Residual Networks (ResNet) models with different sizes and state-of-the-art training recipes from Wightman et al. (2021) and Touvron et al. (2022) (see Appendix A.3 for details). When generating taggants, the signing sets’ sizes use a budget $B$ of the total dataset size, $S=B\times N$, with $B=10^{-3}$ unless stated otherwise. We generate $20$ secret keys, using 8 random restart per key and keep the $K=10$ keys reaching the best taggant objective value. The weight of the perceptual loss is fixed to $\lambda=0.01$. These parameters were chosen to produce visually imperceptible data taggants. $\varepsilon$ in the constraint set $\mathcal{C}$ is fixed to $16/255$, a common value in the data poisoning literature.
In each experiment, we train one model for Alice, craft data taggants, and train Bob’s model on the now protected dataset. We run the detection test with $k=10$ and $\tau=0.1$ (for an associated theoretical FPR of $0.4\%$) and repeat each experiment 4 times with random initializations to compute standard deviations. Similarly to Sablayrolles et al. (2020), we combine the $p$-values of the 4 tests with Fisher’s method (Fisher, 1970). The result is denoted $\log_{10}p$ and is commensurable to the base-$10$ logarithm of a $p$-value. Given the large number of experiments, we train all models for 100 epochs only, while a complete DeIT training is performed on 800 epochs. As sanity check, we confirm with a run of 800 epochs in Table 11 in Appendix A.2 the effectiveness of our method.
Table 1: Comparison of data taggants with DOV baselines when Alice and Bob train a DeIT-small on ImageNet1k with the Three Augment recipe with a budget of $0.1\%$. Aggregated over 4 runs, bold numbers indicate validation accuracy on par with clean training or better, and effective detection.
Method Validation accuracy TPR FPR $\log_{10}p$
Clean training $\mathbf{64.2\pm 0.4}$ - - -
BW – Sleeper Agent
$\mathbf{64.4\pm 0.3}$ $\pm 0.0 0.0±0.0\mathbf$$\pm 0.0$} $(0.0)$
BW – BadNets
$\mathbf{63.7\pm 0.5}$ $\pm 0.0 0.0±0.0\mathbf$$\pm 0.0$} $(0.0)$
Data isotopes $63.0\pm 0.8$ $0.53\pm 0.09$ $0.20\pm 0.08$ -
Data taggants
(Our method)
$\mathbf{64.2\pm 0.6}$ $\mathbf{1.0\pm 0.0}$ $\mathbf{0.0\pm 0.0}$ $\mathbf{(-59.6)}$
Table 2: Validation accuracy and detection performance when both Alice and Bob use DeIT-small models with the three-augment data augmentation with a budget of $0.1\%$.
Method
Validation
accuracy
Top-$k$ keys accuracy PSNR
@@
Naive Canary $63.8\pm 1.1$ $85.0\pm 19.1$ (-91.7) $\mathbf{100.0\pm 0.0}$ (-74.0) -
Transparency $63.6\pm 0.6$ $10.0\pm 0.0$ (-4.9) $\mathbf{55.0\pm 5.8}$ (-29.7) $20.0$
Data taggants
(Our method)
$\mathbf{64.2\pm 0.6}$ $10.0\pm 0.0$ (-4.9) $\mathbf{87.5\pm 5.0}$ (-59.6) $\mathbf{27.9}$
@@
Key
source
Budget
$B$
Validation
accuracy
Top-$k$ keys accuracy PSNR
@@
None $0.0\%$ $64.2\pm 0.4$ - - - - -
Test images
$0.001\%$ $63.7\pm 0.4$ $0.0\pm 0.0$ (0.0) $7.5\pm 5.0$ (-1.1) $27.5$
$0.01\%$ $63.4\pm 1.0$ $0.0\pm 0.0$ (0.0) $0.0\pm 0.0$ ( 0.0) $27.9$
$0.1\%$ $63.9\pm 0.6$ $0.0\pm 0.0$ (0.0) $27.5\pm 5.0$ (-10.4) $28.4$
Random
(Our method)
$0.001\%$ $63.7\pm 0.9$ $0.0\pm 0.0$ (0.0) $0.0\pm 0.0$ ( 0.0) $26.6$
$0.01\%$ $63.6\pm 0.9$ $0.0\pm 0.0$ (0.0) $5.0\pm 5.8$ (-0.5) $27.3$
@@ 4.1 Effectiveness
We first compare the effectiveness of data taggants to three baselines (Section A.1 in Appendix explain this choice) for dataset ownership verification. First, Backdoor watermarking (BW) (Li et al., 2023), rely on a backdoor attack to embed a backdoor behavior in the model. To perform the backdoor attack, we leverage Sleeper Agent (Souri et al., 2021), an effective clean-label backdoor attack, and BadNets Gu et al. (2019), which applies a trigger on images from a source class and flips their labels to a target class. Finally, Data isotopes (Wenger et al., 2022), unlike BW, do not try to induce a backdoor behavior $f(x)=y\neq f(x+x^{(trigger)})$, but only to induce a slight change in the predicted logits. In this comparison, we use DeIT-small models with the Three Augment (3A) data augmentation and associated training recipe (Touvron et al., 2022) for Alice and Bob. We report the validation accuracies, the measured TPR and FPR ($\in[0,1]$), and the $\log_{10}p$-value in Table 1.
Our experiments reveal that in a practical setting, BW is ineffective and yields a $0.0$ detection rate, and data isotopes offer a low detection rate of $0.53$ with a prohibitively high FPR of $0.20$ with a high toll on the validation accuracy. In contrast, data taggants achieve a perfect detection with a $1.0$ TPR and $0.0$ FPR, and a very high confidence of $p<10^{-59}$, without loss of performance.
To measure the effectiveness of gradient matching to craft data taggants and force Bob’s model to learn the keys, we compare it with two baselines to achieve the same goal. First, “naive canary”, where copies of the private keys are added into the training set. This serves as a topline in terms of detection performance, but is not viable as DOV mechanism due to its lack of stealthiness. Second, “transparency”, where we linearly interpolate between the keys and images of the signing set with a weight $\gamma=0.2$, $x^{\prime}=\gamma x^{(key)}+(1-\gamma)x$. This value for $\gamma$ was chosen by visual inspection, as a small value for which the key is still visible. In this comparison, we use DeIT-small models with the three-augment (3A) data augmentation (Touvron et al., 2022) for Alice and Bob, with three different budgets $B$, $1\%, 0.0$$01\%$ and $0.001\%$.
The results are given in Table 2 (more results in Table 7 in Appendix). First, we observe that all methods have roughly the same validation accuracy as a model trained on clean data (differences within error). In particular, this suggests that data taggants do not negatively impact model performances at these budget levels. In terms of detection, Naive canary works best as expected, with perfect key top-10 accuracy for $B\geq 0.01\%$. While Transparency works better than data taggants for smaller budgets, data taggants has much higher top-10 accuracy for $B=0.1\%$ ($87.5\%$ vs $55.0\%$). In addition, the PSNR (signal to noise ratio) is lower for transparency, which suggests that with a budget as small as $0.1\%$, data taggants already dominate the transparency baseline in this experimental setting, and can detect dishonest models with very high confidence ($\log_{10}p=-59.6$).
We perform an additional comparison with another baseline where test data points are used as keys rather than random patterns, which is similar to repurposing a data poisoning attack for DOV. Table 3 shows the results. With comparable PSNR, using random keys leads to much better detection accuracy and confidence for $B=0.1\%$ than using test data, justifying our design choice.
@@ 4.2 Stealthiness
The stealthiness of data taggants is essential but difficult to measure beyond PSNR. As a best effort, we address the scenario where Bob tries to locate the signed data using either visual inspection, defense mechanisms against data poisoning, or, following (Tang et al., 2023), using anomaly detection algorithms. Similarly to the previous section, we use taggants crafted by DeIT-small using 3A augmentation, with a budget $B=0.1\%$.
@@
We provide examples of taggants crafted with and without perceptual loss in Figure 3 (more examples are given in Figure 6, 7 and 9 in Appendix A.2). We observe that while most of them appear unaltered, some data taggants display visible patterns of weak intensity that could as well come from natural film grain or compression artifacts. Although with PSNR << 30 dB (see Table 2) which is considered low according to image processing standards (lower than e.g. Sablayrolles et al. (2020)), we believe that the data taggants are hardly detectable via visual inspection, particularly when Bob has to find them among a whole dataset.
@@
Defense against data poisoning.
Since our method relies on a data poisoning mechanism, we suggest to use data poisoning detection methods to detect data taggants. We consider three data poisoning detection methods relying on filtering samples: Deep k-NN (Peri et al., 2020), Activation Clustering (Chen et al., 2018) and Spectral Clustering (Tran et al., 2018). Over a wide range of parameters, the Receiver Operating Characteristic (ROC) curves in Figure 5 shows that data taggants cannot be detected by these methods with much better performances than random guess, suggesting strong stealthiness of data taggants.
@@
Most methods for out-of-distribution (OOD) detection (Chen et al., 2020) rely on training a model on a set of in-lier data to then be able to detect outliers. In reality Bob does not have access to the original clean data nor to a benign model, hence cannot tell the outliers apart.
We consider the DBSCAN clustering method (Schubert et al., 2017) from the scikit-learn project [^2] https://scikit-learn.org/stable/index.html and run it with a wide range of thresholds ($\in[10,35]$) to cluster the features of a model $h$ trained on $\tilde{\mathcal{D}}_{A}$. The resulting ROC curves are given in Figure 5 as a scatter plot after the experiment 4 times. Interestingly, when computing DBSCAN for different detection thresholds, we observe that it exhibit performances that are significantly worse than random. Our explanation is that DBSCAN computes clusters and select outliers as isolated points or clusters. By manually analyzing the clusters, we observe that some of them contain a lot of data taggants, most likely because their embeddings are lumped together and do not appear as anomalies.
@@
Figure 4: ROC curves for defense against data poisoning methods in the detection of data taggants.
Figure 5: ROC curve for DBSCAN anomaly detection method.